the sum of the ages of Petra and her mother is 53. her mother is 5 years more than twice as old as Petra. How old are petra and her mother?
Answer by JulietG(1812)

Let P be Petra's age.
Let M be her mother's age.
Here's what we know from the question:
P + M = 53
M = 2P + 5
We can substitute the value of M into the first equation:
P + (2P+5) = 53
Add
3P + 5 = 53
Subtract 5 from each side:
3P = 48
Divide each side by 3:
P = 16
.
If Petra is 16, then her mother must be 37 (53-16=37)